OpenStudy (anonymous):

sinx/1+cosx = cscx-cotx prove ls=rs

OpenStudy (anonymous):

can you help me with something? can you give me a little more detail about the probelm/

OpenStudy (kainui):

I always hated this kind of useless question.

OpenStudy (anonymous):

same here.. I mean it is not easy but if you know your stuff you solve it.

OpenStudy (kainui):

I'm afraid I don't. I was thinking it looked like you could turn it into sinx+tanx but really I'd rather do a trig substitution integral any day, good luck! =P

OpenStudy (anonymous):

well how about we work backwards to get to the answer

OpenStudy (anonymous):

\[sinx \div 1+cosx = cscx - cotx\]

OpenStudy (anonymous):

let us start with cscx-cotx tell me what that is?

OpenStudy (anonymous):

I don't think this question is right . . .

OpenStudy (anonymous):

question is correct.

OpenStudy (anonymous):

to do this problem use reciprocal functions

OpenStudy (anonymous):

write the right side as \[\Large \frac{1}{\sin(x)}-\frac{\cos(x)}{\sin(x)}\] now take LCM sin(x) \[\Large \frac{1-\cos(x)}{\sin(x)}\] now multiply and divide the above with 1+cos(x) you will get the left side.

OpenStudy (anonymous):

hey same 21 just wondering you could do that backwards as well and get to the same answer?

OpenStudy (anonymous):

I think i'd lile to see that as well.

OpenStudy (anonymous):

wait how do you get cosx/sinx from cotx?

OpenStudy (anonymous):

@godorovg you mean to go from left side to prove right side ??

OpenStudy (anonymous):

yes that is what I mean

OpenStudy (anonymous):

always. \[\Large \cot(x)=\frac{\cos(x)}{\sin(x)}\] since \[\Large \tan(x)=\frac{\sin(x)}{\cos(x)}\] and \[\Large \cot(x)=\frac{1}{\tan(x)}\]

OpenStudy (anonymous):

oh okayy

OpenStudy (anonymous):

see it works both ways sami-21 or I am wrong?

OpenStudy (anonymous):

it works both ways.

OpenStudy (anonymous):

sami- 21 i find it easier to work backwards with these problems it seems easier too me. you?

OpenStudy (anonymous):

i find it easier the other way. in backwards approach you will have to do more steps.

OpenStudy (anonymous):

sami -21 I know in some cases backwards won't work.